OpenStudy (christos):

can someone tell me how the notation for the double derivative , d^2y/dx^2 is derived from dy/dx ?

Parth (parthkohli):

I read an explanation that\[\frac{d}{dx}\left(\frac{dy}{dx}\right) = \frac{d^2y}{dx^2}\]but you shouldn't take it too seriously

OpenStudy (christos):

hm

OpenStudy (mathmate):

Leibniz's notation. see: https://en.wikipedia.org/wiki/Notation_for_differentiation